EVA Report – March 4th

Crew 206 EVA Report 04-MAR-2019
EVA #5

Author: Gaspard Thieulin

Purpose of EVA:

The main purpose of the EVA was to do the basic checks of our new ‘SUPAERO SpaceSuit’. We brought a prototype of a MDRS suit which worked on it for a year. On wednesday we worked inside the hab on the protocol for pre-EVA and post EVA manipulations. Everything was nominal. Now we wanted to try it just in front of the hab.

Here is what is planned :
– Ability to realize an EVA :
– Isolation from the outside environment
– Providing air to the astronaut
– Providing water to the astronaut
– Enabaling mobility (arms, back and neck) for the astronaut – Ability to communicate with the crew
– Providing a good vision through the helmet
– Providing radio communication
– Protecting the astronaut from the sun
– Deploying the sun protection easily

Once those points were checked, we had to try this second protocole :

– Ability to remove the fog from the helmet
– Try to seat in the rovers without moving
– Try to simulate drivng conditions

But first, the team “Red” had to quickly go to the LOAC to change the battery, and check the weather station orientation after this morning’s earthquake.

Finally, the EVA teams were also meant to fix the tarp of the corridor between the science dome and the Hab.

Start time: 14h19

End time: 15h26

Narrative: Today’s EVA was conditionnaly approved, initially for the morning. We had to ask for MS’ approval before going out. In the morning, Atila went outside and experienced a very muddy ground and refused the EVA for the morning. However, he decided to let us decide if we wanted to conduct the EVA in the afternoon. The crew decided to take a decision by 1400 for an EVA replanned on 1415.

Having this decision in our hand was a lot of responsibility, especially for the EVA leader. No matter what we would decide in the afternoon, we decided to review a bit our EVA plan and add the maximum of Safety to it . Only two crew members (team Red: Aurelien and myself as EVA leader) would go outside first, taking one final decison on the pursuing of the EVA and looking for a dry area to conduct the prototype tests. Only then they would go for engineering checks and LOAC maintenance. During this whole time, the two other crew members (team Blue: Benjamin and Norbert who would try the prototype suit) would wait in the Hab, wearing EVA suits to be able to save the others as quick as possible.

Only after team Red comes back to the Hab area, one member of team Blue would swap his EVA suit for the prototype one, and go outside to conduct the rest of the EVA.

EVA leader would still be able to cancel the EVA by any time, of course.

We began equipping ourselves at 13h30 to be able to go out at 14h15, if we decided to go out. At 1400 today we decided to take a shot at this EVA’s plan, considering the whole day had been sunny and rather “hot”. Mission support (Atila in this case) accepted our plan and decision.

We went out of the airlock at 14h19. My first step on the ground would be decisive, and we were extremly pleased to discover that the ground had almost completly dried out during the day. In fact, it was very easy to move on dry dust, and the rare muddy zones were visible enough not to represent any danger. Considering that, I decided to conduct the EVA as planned and we successfully conducting every check, changed the LOAC battery and drived back while team Blue was waiting in the hab fully suited. We parked Curiosity at 14h35 and team Blue began
depressurization while Norbert was wearing the prototype suit. They also brought back some fixations for the tarps in the prototype backpack’s storage.

We were then able to conduct every test in the EVA’s protocol, and gathered a lot of data thanks to it. We ended these tests at 15h12, and Norbert was already warning us about his stamina decreasing fast due to the weight of the prototype suit.
We still were able to fix some parts of the teared tarps with the material brought in the suit in a very limited time. The fixations we used are lighter than the previous one, and we didn’t have enough to attach every part of the tarp, so we plan on another fixing session soon. Finally, we came back in the airlock at 15h21 and the EVA successfully ended at 15h26.

Destination: Hab area and LOAC / weather station.

Coordinates: 518500/425700 (LOAC and Weather Station) ; 518150/4250750 (Hab)

EVA Participants: Benjamin Auzou (COMMS), Gaspard Thieulin (ENG, Eva Leader), Aurélien Mure (XO), Norbert Pouzin (GHO, prototype Suit)

Roads and routs per MDRS Map: Cow Dung Road

Mode of travel: Walking and Rover (for carrying our heavy battery)

Vehicles used : Curiosity

Copyright © The Mars Society. All rights reserved. | Main Site